# UniCon: A Unified System for Efficient Robot Learning Transfers

### v1

@@ 1 Introduction
To address above challenges, we present UniCon, a lightweight framework for robot learning deployments that provides the unifying infrastructure needed to standardize states, control flow, and instrumentation across platforms. UniCon decomposes workflows into execution graphs with reusable basic units, and separates system states – such as proprioceptive, exteroceptive, and control signals – from the logic that consumes them. This enables plug‑and‑play deployment across diverse morphologies including quadrupeds, humanoids, and manipulators, while unifying hardware with simulators for seamless transfer.
Compared to ROS, our architecture adopt a data-oriented modular paradigm with separation between runtime states and control logic. To maximize efficiency, we favor single-threaded computing (ros2compose) over multiprocessing, and vectorized data flow over message passing. On the other hand, interoperability with existing software is still guaranteed by extending the underlying states storage to multiple communication backends (zeromq2010; ddsfoundation2025). We show that this reduces code redundancy and improves inference latency in our cross-embodiment experiments.
@@
UniCon has already been deployed on over 12 robot models spanning 7 manufacturers, and has supported 3 research projects ongoing in the Shanghai Artificial Intelligence Laboratory, demonstrating its versatility in real research scenarios. We release our framework as open-source at https://github.com/creeperlin/unicon.
@@ 2 Framework
A key obstacle in sim‑to‑real and cross-embodiment transfer is the tight coupling of system data flows with customized control logic. Middleware such as ROS enforce a message passing pattern but incurs latency, making it unsuitable for learning‑based controllers. Our design instead adopts a modular, data‑oriented paradigm that cleanly separates states from logic, emphasizing reusability and efficiency. This approach contrasts with object‑oriented programming practices (10.5555/1407387), where data and routines are encapsulated together, and instead aligns more closely with the Entity Component System architecture widely used in the game industry (overwatch2017gdc) and modern simulation engines (shacklett23madrona).
@@ 2.1 Vectorized States
We formulate all runtime states of a robot workflow as global array objects of numerical data types under different labels. This comes from the fact that the data passed in structured messages often take a fixed format for each channel if the hardware participants remains constant. Therefore, rather than transmitting full self‑describing objects, only the essential multidimensional data are exchanged, eliminating unnecessary encapsulation and aligning with contiguous memory layouts favored by simulators and hardware interfaces.
Typically, the states includes robot sensory inputs such as motor position and velocities, as well as orientation and angular velocity reported by the IMU. Similarly, motor commands such as position and torque controls are also states which are written to by controllers and read from by hardware interfaces. In the case where state definitions of custom components differ from the actual hardware, transparent mappings with efficient vectorized indexing operations is inserted to align properties such as joint order.
@@ 2.2 Modularized Logic
A workflow is then constructed by nesting the CBs, forming an execution graph. We provide basic control flow components in similar fasion to functional programming (Hughes1989WhyFP) for composing customized workflows. For example, loop terminates the inner block when the predicate is satisfied, zip executes multiple blocks in sequence, and chain exhausts nested blocks in order until they all halts.
@@
This enables configurations via text formats and possibly GUIs, while also leaves directives for code generation.
@@ 2.3 Software Interoperability
Peripherals.  We also incoporate frequently used peripheral tools and devices as plug-and-play CBs. Input devices such as joysticks (linuxinput2025) and VR (vuer) are supported with multiple implementations for coverage, while sensors including RGBD (realsense) cameras and LiDARs are connected either through platform proxy or standalone clients. Visualizers such as Foxglove and Meshcat (qin2023anyteleop) are also ready to be wired to selected states.
@@ 2.5 Workflow Transfers
UniCon enables seamless tranfer between three workflows: Sim-to-Sim validation, Sim-to-Real deployment, and Real-to-Sim instrumentation, all operating on a unified data abstraction. Sim-to-Sim and Sim-to-Real are trivial since the system CBs remain interchangeable within the execution graph. For real deployments, additional safety blocks can be inserted to handle startup, shutdown, and runtime monitoring without altering the rest of the graph.
Real-to-Sim uses recorder and replay blocks on physical states and built‑in metrics to quantify the reality gap. For open‑loop and repeatable rollouts, mean squared error (MSE) performs step‑wise comparison since inputs are exactly replayed. For closed‑loop controls like locomotion, we propose to unfold the MSE loss to prevent misalignments:
$\mathcal{L}=\frac{1}{n}\min_{j}\sum_{i=0}^{n}\|T_{i+j}-\hat{T}_{i}\|_{2}+\frac{1}{n}\min_{j}\sum_{i=0}^{n}\|T_{i}-\hat{T}_{i+j}\|_{2}.$ (4)
@@ 3 Evaluation
Gazebo (Koenig2004DesignAU) AzureLoong (oghr2)
Newton (newton2025github) Booster T1 (t1)
@@ 3.1 Transfer Effort
We target the locomotion task for humanoids and quadrupeds where policies with the same training paradigm are deployed on multiple platforms (xue2025unified). Without a framework, the inference often need to be re‑implemented due to vendor‑specific stacks. We report the Source Lines of Code (SLOC) required to transfer the original workflow on Unitree H1 to the MuJoCo simulator and other robots. As shown in Table 2, the Ad‑hoc code needs substantial efforts to transfer, especially for the PND Adam model with its completely different APIs, while UniCon remains lightweight and requires no extra effort when switching deployment targets.
@@ 3.2 Framework Efficiency
Results in Table 3 show that UniCon introduces near‑zero overhead compared to SDK implementations and significantly outperforms ROS. This efficiency arises from binding global state buffers into the data reader/writer and compiling (pybind11github; RoboJuDo) alongside the SDK to eliminate redundant data copies. During actual inference, control blocks can be arranged so that the critical path contains only the data producers required by the policy, deferring non‑critical communication and synchronization.
@@ 3.3 Workflow Usability
While ROS is the go-to choice for building complex robotic application, UniCon offers comparable usability with built-in connection to mainstream sensors, input devices, motion libraries and messaging layers including ROS itself. We demonstrate this by extending the inference pipeline to wholebody control, incorporating inverse kinematics (Buss2004IntroductionTI; pink), dexterous hand retargeting (qin2023anyteleop), and VR teleoperation (Cheng2024OpenTeleVisionTW), all modularized and distributed across onboard and LAN workstations.
@@ 4.2 Robot Application Middleware
Robot application middleware provides critical abstractions and communication infrastructure for integrating perception, planning, and control components across heterogeneous robotic systems. Frameworks such as the Robot Operating System (ROS) have become de facto standards in academia and industry due to their modularity, device driver ecosystem, and support for distributed computing (ros; ros2; Chitta2012MoveItT). Similarly, PX4 (Meier2015PX4AN) offers a full-stack flight control system for aerial robotics, facilitating high-performance real-time control with rich simulation integration.
@@ 5 Conclusion
UniCon offers a unified control layer in robot learning deployments by separating vectorized global states from modular control blocks, greatly reducing hardware transfer effort and enabling low-latency inference. Its successful use across robots and research projects show strong potentials for broader community adoption. Future works include multi‑language APIs and code generation to broaden platform coverage.
@@ Acknowledgments
The Shanghai Jiao Tong University team is partially supported by National Natural Science Foundation of China (62322603). We also thank Fourier Intelligence and ByteDance Seed for their early support with hardware and experiments.

### v2

@@ 1 Introduction
To address these challenges, we present UniCon, a lightweight framework for robot learning deployments that provides the unifying infrastructure needed to standardize states, control flow, and instrumentation across platforms. UniCon decomposes workflows into execution graphs with reusable basic units, and separates system states – such as proprioceptive, exteroceptive, and control signals – from the logic that consumes them. This enables plug‑and‑play deployment across diverse morphologies including quadrupeds, humanoids, and manipulators, while unifying hardware with simulators for seamless transfer.
Compared to ROS, our architecture adopts a data-oriented modular paradigm with separation between runtime states and control logic. To maximize efficiency, we favor single-threaded computing (ros2compose) over multiprocessing, and vectorized data flow over message passing. On the other hand, interoperability with existing software is still guaranteed by extending the underlying states storage to multiple communication backends (zeromq2010; ddsfoundation2025). We show that this reduces code redundancy and improves inference latency in our cross-embodiment experiments.
@@
UniCon has already been deployed on over 12 robot models spanning 7 manufacturers, and has supported 3 research projects in the Shanghai Artificial Intelligence Laboratory, demonstrating its versatility in real research scenarios. We release our framework as open-source at https://github.com/creeperlin/unicon.
@@ 2 Framework
A key obstacle in sim‑to‑real and cross-embodiment transfer is the tight coupling of system data flows with customized control logic. Middleware such as ROS enforces a message passing pattern but incurs latency, making it unsuitable for learning‑based controllers. Our design instead adopts a modular, data‑oriented paradigm that cleanly separates states from logic, emphasizing reusability and efficiency. This approach contrasts with object‑oriented programming practices (10.5555/1407387), where data and routines are encapsulated together, and instead aligns more closely with the Entity Component System architecture widely used in the game industry (overwatch2017gdc) and modern simulation engines (shacklett23madrona).
@@ 2.1 Vectorized States
We formulate all runtime states of a robot workflow as global array objects of numerical data types under different labels. This comes from the fact that the data passed in structured messages often takes a fixed format for each channel if the hardware participants remains constant. Therefore, rather than transmitting full self‑describing objects, only the essential multidimensional data are exchanged, eliminating unnecessary encapsulation and aligning with contiguous memory layouts favored by simulators and hardware interfaces.
Typically, the states include robot sensory inputs such as motor position and velocities, as well as orientation and angular velocity reported by the IMU. Similarly, motor commands such as position and torque controls are also states which are written to by controllers and read from by hardware interfaces. In the case where state definitions of custom components differ from the actual hardware, transparent mappings with efficient vectorized indexing operations are inserted to align properties such as joint order.
@@ 2.2 Modularized Logic
A workflow is then constructed by nesting the CBs, forming an execution graph. We provide basic control flow components in similar fashion to functional programming (Hughes1989WhyFP) for composing customized workflows. For example, loop terminates the inner block when the predicate is satisfied, zip executes multiple blocks in sequence, and chain exhausts nested blocks in order until they all halt.
@@
This enables configurations via text formats and possibly GUIs, while also leaving directives for code generation.
@@ 2.3 Software Interoperability
Peripherals.  We also incorporate frequently used peripheral tools and devices as plug-and-play CBs. Input devices such as joysticks (linuxinput2025) and VR (vuer) are supported with multiple implementations for coverage, while sensors including RGBD (realsense) cameras and LiDARs are connected either through platform proxy or standalone clients; high-bandwidth image data are handled through their native interfaces and referenced via lightweight handles to preserve efficiency. Visualizers such as Foxglove and Meshcat (qin2023anyteleop) are also ready to be wired to selected states.
@@ 2.5 Workflow Transfers
UniCon enables seamless transfer between three workflows: Sim-to-Sim validation, Sim-to-Real deployment, and Real-to-Sim instrumentation, all operating on a unified data abstraction. Sim-to-Sim and Sim-to-Real are trivial since the system CBs remain interchangeable within the execution graph. For real deployments, additional safety blocks can be inserted to handle startup, shutdown, and runtime monitoring without altering the rest of the graph.
Real-to-Sim uses recorder and replay blocks on physical states and built‑in metrics to quantify the reality gap. For open‑loop and repeatable rollouts, mean squared error (MSE) performs step‑wise comparison since inputs are exactly replayed. For closed‑loop controls like locomotion, we propose to unfold the MSE loss to prevent temporal misalignments:
$\mathcal{L}=\frac{1}{n}\min_{j}\sum_{i=0}^{n}\|T_{i+j}-\hat{T}_{i}\|_{2}+\frac{1}{n}\min_{j}\sum_{i=0}^{n}\|T_{i}-\hat{T}_{i+j}\|_{2},$ (4)
where $T$ represents the real robot trajectory, $\hat{T}$ represents the simulated trajectory, and $j$ is the temporal shift minimized to find the best alignment between the two sequences.
@@ 3 Evaluation
Gazebo (Koenig2004DesignAU) AzureLoong (oghr2) Inspire RH56
Newton (newton2025github) Booster T1 (t1) XHand1
@@ 3.1 Transfer Effort
We target the locomotion task for humanoids and quadrupeds where policies with the same training paradigm are deployed on multiple platforms (xue2025unified). Without a framework, the inference often needs to be re‑implemented due to vendor‑specific stacks. We report the Source Lines of Code (SLOC) required to transfer the original workflow on Unitree H1 to the MuJoCo simulator and other robots. As shown in Table 2, the Ad‑hoc code needs substantial efforts to transfer, especially for the PND Adam model with its completely different APIs, while UniCon remains lightweight and requires no extra effort when switching deployment targets.
@@ 3.2 Framework Efficiency
Results in Table 3 show that UniCon introduces near‑zero overhead compared to Python SDK implementations and significantly outperforms ROS. This efficiency arises from binding global state buffers into the data reader/writer and compiling (pybind11github; RoboJuDo) alongside the SDK to eliminate redundant data copies. During actual inference, control blocks can be arranged so that the critical path contains only the data producers required by the policy, deferring non‑critical communication and synchronization.
@@ 3.3 Workflow Usability
While ROS is the go-to choice for building complex robotic applications, UniCon offers comparable usability with built-in connection to mainstream sensors, input devices, motion libraries and messaging layers including ROS itself. We demonstrate this by extending the inference pipeline to wholebody control, incorporating inverse kinematics (Buss2004IntroductionTI; pink), dexterous hand retargeting (qin2023anyteleop), and VR teleoperation (Cheng2024OpenTeleVisionTW), all modularized and distributed across onboard and LAN workstations.
@@ 4.2 Robot Application Middleware
Robot application middleware provides critical abstractions and communication infrastructure for integrating perception, planning, and control components across heterogeneous robotic systems (Gerkey2003ThePP; Metta2006YARPYA; Bruyninckx2001OpenRC; Utz2002MiroM). Frameworks such as the Robot Operating System (ROS) have become de facto standards in academia and industry due to their modularity, device driver ecosystem, and support for distributed computing (ros; ros2; Chitta2012MoveItT). Similarly, PX4 (Meier2015PX4AN) offers a full-stack flight control system for aerial robotics, facilitating high-performance real-time control with rich simulation integration.
@@ 5 Conclusion
UniCon offers a unified control layer in robot learning deployments by separating vectorized global states from modular control blocks, greatly reducing hardware transfer effort and enabling low-latency inference. Its successful use across robots and research projects shows strong potential for broader community adoption. Future work includes multi‑language APIs and code generation to broaden platform coverage.
@@ Acknowledgments
The SJTU team is supported by National Natural Science Foundation of China (62322603), Shanghai Municipal Science and Technology Major Project (2025SHZDZX025D08) and Shanghai Artificial Intelligence Laboratory. We also thank Fourier Intelligence and ByteDance Seed for their early support with hardware and experiments.